# On the geometry of punctual Hilbert schemes on singular curves and their motivic zeta functions

### v3

@@ Abstract.
Inspired by the work of Soma and Watari, we define a tree structure on certain subsemimodules of the semigroup $\Gamma$ associated with an irreducible plane curve singularity $(C,O)$. Building on results of Oblomkov, Rasmussen, and Shende, we show that for specific classes of singularities, this tree encodes key aspects of the geometry of the punctual Hilbert schemes of $(C,O)$. As an application, we compute the motivic Hilbert zeta function for a family of singular curves.
A point in the Hilbert scheme corresponds to an ideal in the local ring $\mathcal{O}_{C,O}$ of the singularity. We study the stratification of these Hilbert schemes induced by constraints on the minimal number of generators of the defining ideals, and we describe geometric properties of these strata, including their dimension and closure relations.
More importantly, we study their motivic zeta functions, particularly the motivic Hilbert zeta function, which encodes the classes of all punctual Hilbert schemes in the Grothendieck ring of varieties.
@@ 1. Introduction
From an algebraic viewpoint, let $A=\mathcal{O}_{(C,O)}$ denote the local ring of $C$ at $O$. Then the $k$-points of $C^{[\ell]}$ are in bijection with ideals of colength $\ell$ in $A$:
@@
Pfister and Steenbrink [16] proved that the punctual Hilbert schemes $C^{[l]}$ can be embedded in a closed linear subvariety of the Grassmannian $Gr(\delta,\overline{A}/I(2\delta)),$ where $I(2\delta)$ is the ideal consisting of elements $h\in\overline{A}$ (or equivalently of $A$) whose valuation $v(h)$ is larger than or equal to $2\delta.$ Moreover, the authors ”provided” the defining equations of this embedding. In general, even when the defining equations of punctual Hilbert schemes are known, it remains difficult to understand their geometry or to compute their invariants.
Progress in understanding the geometry of punctual Hilbert schemes has been relatively slow, due in part to the complexity of their structure even in simple singular cases. Soma and Watari [21, 22] studied the geometry of punctual Hilbert schemes for plane curve singularities of types $A_{2d}$, $E_{6}$, and $E_{8}$. In the case of unibranched curve singularities, the punctual Hilbert scheme $C^{[\ell]}$ coincides with the compactified Jacobian $\overline{JC}$ for sufficiently large $\ell$ [17].
For curves defined by the equation $y^{m}-x^{n}=0$ with $\gcd(m,n)=1$, Lusztig and Smelt [12] computed the homology of the associated Hilbert schemes. Piontkowski [17] later extended these results to curves with a single Puiseux pair. Subsequently, Gorsky, Mazin, and Oblomkov [5] derived explicit formulas for the Poincaré polynomials of Hilbert schemes associated with singularities having Puiseux exponents $(nd,md,md+1)$.
A key insight in this line of research is that the intersection of the compactified Jacobian $\overline{JC}$ with certain Schubert cells in the affine Grassmannian has the structure of an affine space. This geometric fact admits rich combinatorial interpretations, particularly in relation to $(q,t)$-Catalan theory [6, 7, 8, 10].
For curve singularities with a single Puiseux pair, Oblomkov, Rasmussen, and Shende [14] proved that the punctual Hilbert scheme $C^{[\ell]}$ admits a stratification by affine spaces, a result that plays a foundational role in the motivic study of these moduli spaces.
@@
On one hand, these schemes are the fundamental building blocks for the motivic Hilbert zeta function, introduced by Bejleri, Ranganathan and Vakil [1]. This zeta function is defined as the generating series of the classes of punctual Hilbert schemes in the Grothendieck ring of varieties:
@@
Compared to the classical motivic zeta function of Kapranov [11], the motivic Hilbert zeta function provides a finer invariant, capturing detailed information about the singularities of a variety $X$ defined over an algebraically closed field $k$.
On the other hand, a series of conjectures by Oblomkov, Rasmussen, and Shende [15, 14] posits a deep connection between the geometry of punctual Hilbert schemes of plane curve singularities and the topology of associated algebraic knots, particularly over $\mathbb{C}$. More precisely, they conjecture that the (reduced) Khovanov–Rozansky homology of the link of the singularity can be recovered from the homology of the Hilbert schemes. Maulik [13] proved the first of these conjectures using an inductive argument based on blow-ups, analyzing the variation of both Hilbert schemes and the HOMFLY polynomial under such transformations. Notably, his approach yields significant geometric insight into the structure of punctual Hilbert schemes.
Very recently, Rossinelli [19] discovered a new connection between curvilinear punctual Hilbert schemes and the Igusa motivic zeta function, further highlighting the role of Hilbert schemes as a unifying object in singularity theory and motivic integration. Diaconescu, Porta, Sala and Vosoughinia [3] made a progress in the case of space curve singularities. Fasola, Graffeo, Lewanski and Ricolfi [4] computed the motivic classes of punctual nested Hilbert schemes on a smooth pointed surface. Graffeo, Monavari, Moschetti, and Ricolfi [9] proved a closed formula for the generating function of the motives $\mathrm{Hilb}^{d}(\mathbb{A}^{n})_{0}]\in K_{0}(\mathrm{Var}_{\mathbb{C}})$ of punctual Hilbert schemes, summing over $n$, for fixed $d>0$.
@@
The latter class consists of singularities for which, within their equisingularity class—i.e., the set of curve singularities sharing the same semigroup—there exists a unique representative up to analytic isomorphism (see Section 6 for futher details). This class was originally introduced by Pfister and Steenbrink [16], and plays a distinguished role in the classification of unibranch curve singularities due to its rigidity under deformation.
To achieve this, we make extensive use of $\Gamma$-subsemimodules. A $\Gamma$-subsemimodule is a subset $\Delta\subset\Gamma$ such that $\Gamma+\Delta\subseteq\Delta$. It follows directly from the properties of ideals and the valuation $v$ that, for any ideal $I\subset A$, the value set
@@
Inspired by recent work of Soma and Watari [22] on the geometry of Hilbert schemes, as well as that of Oblomkov, Rasmussen, and Shende [14] on knot invariants and curve singularities, we associate to a curve singularity with semigroup $\Gamma$ a leveled graph $G_{\Gamma}$ defined as follows.
@@
We study the graph $G_{\Gamma}$ and prove that it has the structure of a tree (Theorem 3.1). Furtheromore, each subsemimodule $\Delta\in\mathscr{D}_{\ell}$ defines a constructible subset $C^{[\Delta]}\subseteq C^{[\ell]}.$ One of the main result of the paper, building on the work in [14] regarding the defining equations of $C^{[\Delta]}$, is that for certain curve singularities an edge of the graph induces a peicewise trivial fibration:
We study the graph $G_{\Gamma}$ and prove that it has the structure of a tree (Theorem 3.1). Furthermore, each subsemimodule $\Delta\in\mathscr{D}_{\ell}$ defines a constructible subset $C^{[\Delta]}\subseteq C^{[\ell]}$, consisting of all ideals $I\subset A$ in $C^{[\ell]}$ such that $v(I\setminus\{0\})=\Delta$.
One of the main results of this paper, building on the description of defining equations for $C^{[\Delta]}$ due to Oblomkov, Rasmussen, and Shende [14], is that for certain curve singularities—particularly those with monomial semigroups—an edge in the tree $G_{\Gamma}$ induces a piecewise trivial fibration. More precisely, we prove the following.
@@
We explicitly determine the image of this morphism. In the case of plane curve singularities defined by $x^{p}-y^{q}=0$ with $\gcd(p,q)=1$, or more generally for singularities with monomial semigroups, this allows us to recover a key result from [14]: $C^{[\Delta]}$ is an affine space, and its dimension is completely determined by the combinatorial invariants of the $\Gamma$-subsemimodule $\Delta$.
We then analyze an example, see Example 4.3, of a plane curve singularity with two Puiseux pairs. In this case, the morphism
@@
still induces a piecewise trivial fibration, but the dimension of the fiber is not solely governed by the number of minimal generators below $\gamma_{\Delta}$. New phenomenas arise and must be further studied to fully understand the geometry of the strata and the structure of the fibration.
One important application of Theorem 1.1 is the development of an algorithm to compute the motivic Hilbert zeta function for irreducible curve singularities, as introduced in [1]:
@@
For the singularities of types $A_{2d}$, $E_{6}$, and $E_{8}$, these formulas were independently obtained by Watari [23], using a different approach based on explicit geometric analysis and results from [22, 21]. Our method, based on the tree structure $G_{\Gamma}$ and piecewise trivial fibrations, provides a unified and algorithmic derivation that extends naturally to more general singularities, including $W_{8}$ and $Z_{10}$.
In the Section 5 , in relation to the conjectures of Oblomkov, Rasmussen, Shende, and Maulik [15, 14, 13], we study the defining equations of certain stratifications of the punctual Hilbert schemes. For a given $\Gamma$-subsemimodule $\Delta$ and an integer $m\in\mathbb{N}$, we consider the
@@
where $n$ is the number of minimal generators of $\Delta$ as a $\Gamma$-subsemimodule, and $N(\Delta)$ is the dimension of $C^{[\Delta]}$ which is an affine space proven in [14, Theorem 13].
@@
This series $Z^{\mathrm{Hilb}}_{(C,O)}(a^{2},q^{2})$, introduced in this paper, generalizes the motivic zeta function studied in [15]. There, the Euler characteristic specialization (i.e., the point-counting or topological Euler characteristic) of this series was conjecturally related to the HOMFLY polynomial of the associated algebraic knot. Our construction lifts this invariant to the Grothendieck ring of varieties $K_{0}({\mathrm{Var}}_{{\mathbb{C}}})$, thereby encoding richer geometric information about the stratification of the Hilbert schemes by the number of generators.
@@ 2. Punctual Hilbert schemes
In this section, we recall the definition of the punctual Hilbert schemes of a curve singularity and review some fundamental properties. Let $(C,O)$ be the germ of a unibranch curve singularity defined over an algebraically closed field $k$ of characteristic zero, and let $A:=\mathcal{O}_{C,O}$ denote its local ring at $O$. Then $A$ is a complete one-dimensional local domain, and its integral closure $\overline{A}$ in its field of fractions is a discrete valuation ring, isomorphic to $k[[t]]$.
@@
and $c=2\delta$ if and only if $A$ is Gorenstein [20].
@@
Pfister and Steenbrink [16] showed that the punctual Hilbert schemes of a unibranch curve singularity can be embedded as closed subvarieties of a Grassmannian. More precisely, let $\mathscr{M}$ denote the reduced subscheme of the Grassmannian $\operatorname{Gr}(\delta,\overline{A}/I(2\delta))$ defined by
@@
Since $c\leq 2\delta$, this implies that for $\ell\geq 2\delta$, the structure of $C^{[\ell]}$ stabilizes in the sense that its points are completely determined by the fixed ambient variety $\mathscr{M}$. Consequently, to understand the classes $C^{[\ell]}]$ in the Grothendieck ring of varieties $K_{0}(\mathrm{Var}_{k})$, it suffices to study the punctual Hilbert schemes for $\ell<2\delta$.
@@
[18, Lemme 5.1.24] For positive integer $\ell$, an ideal $I$ of $A$ belongs to $C^{[\ell]}$ if and only if $v(I\setminus\{0\})\in\mathscr{D}_{\ell}$.
@@
It is shown in [16, Lemma 5] that each stratum $C^{[\Delta]}$ arises as the intersection of $C^{[\ell]}$ with a Schubert cell in the Grassmannian $\operatorname{Gr}(\delta,\overline{A}/I(2\delta))$. In particular, this implies that $C^{[\Delta]}$ is a open subset of $C^{[\ell]}$, and its geometry is compatible with the cellular decomposition of the ambient Grassmannian.
@@ 3. The tree structure of subsemimodules
Let $\Gamma$ be the value semigroup associated with the germ of an irreducible curve singularity $(C,O)$. The goal of this section is to endow the set of $\Gamma$-subsemimodules with a tree structure. We will show in the next section that this combinatorial object plays a crucial role in understanding the geometry of the punctual Hilbert schemes of $C$. This construction is inspired by the work of Soma and Watari [22, 21].
@@
The following result appears in [22]:
@@
The answer is affirmative, as will follow from Proposition 3.1, which shows that the deletion maps are jointly surjective. To prepare for this result, we first establish the following two lemmas.
@@
    Acyclicity: t follows from the uniqueness of Frobenius element of a $\Gamma$-subsemimodule.
@@ 4. Piecewise fibrations induced by the edges of the $\Gamma$-subsemimodules tree
Corollary 4.2 provides a new perspective on the geometry of $C^{[\Delta]}$ in the case $C=\{y^{p}=x^{q}\}$, recovering a result from [14, Theorem 13] via our combinatorial and valuative approach. Our method relies on the explicit defining equations of the strata $C^{[\Delta]}$, introduced in [14, Proposition 12].
We begin by recalling some key constructions from [14]. Let $(C,O)$ be the germ of a unibranch curve singularity—either of type $(p,q)$ or with a monomial semigroup—and let $A:=\mathcal{O}_{C,O}\subset{\mathbb{C}}[[t]]$ denote its complete local ring. We fix a basis of $A$ that is compatible with the monomial basis of the semigroup algebra ${\mathbb{C}}[\Gamma]$. Specifically, for each $i\in\Gamma$, we consider elements of the form
@@
$f_{\gamma_{j}}=\phi_{\gamma_{j}}+\sum_{k\in\Gamma_{>\gamma_{j}}\setminus\Delta}\lambda_{j}^{k-\gamma_{j}}\phi_{k}$
@@
$\tau_{\gamma_{j}}(\lambda^{\bullet}_{\bullet})=\phi_{\gamma_{j}}+\sum_{k\in\Gamma_{>\gamma_{j}}\setminus\Delta}\lambda_{j}^{k-\gamma_{j}}\phi_{k},\quad\lambda_{j}^{k-\gamma_{j}}\in{\mathbb{C}}.$
@@
This bijectivity follows from [15, Theorem 27], which ensures that every ideal in $C^{[\Delta]}$ admits a unique normalized generating set of the prescribed form. Consequently, the stratum $C^{[\Delta]}$ embeds into the parameter space $\operatorname{Gen}_{\Delta}$, yielding a locally closed embedding
@@
It is important to note that the inclusion $C^{[\Delta]}\hookrightarrow\operatorname{Gen}_{\Delta}$ does not imply that every point in $\operatorname{Gen}_{\Delta}$ corresponds to an ideal $J$ with $v(J\setminus\{0\})=\Delta$.
This phenomenon is illustrated by the following example, adapted from [15]. Let $\mathcal{O}_{C}={\mathbb{C}}[[t^{3},t^{7}]]$, whose value semigroup is $\Gamma=\langle 3,7\rangle$. Consider the $\Gamma$-subsemimodule $\Delta=(6,10)_{\Gamma}$, minimally generated by $6$ and $10$. The element
@@
where $S$ denotes the composition ${\mathbb{C}}[\Gamma]^{\oplus m}\to\ker G\hookrightarrow{\mathbb{C}}[\Gamma]^{\oplus n}$. According to [17, Lemma 4], the kernel $\ker G$ is generated by homogeneous syzygies of the form
@@
$(s_{i})_{j}=u_{i}^{j}t^{\sigma_{i}-\gamma_{j}},\quad u_{i}^{j}\in{\mathbb{C}},$
and only finitely many $u_{i}^{j}$ are nonzero.
@@
1.We define the Augmented Syzygy of $\Delta$ as

### v4

@@ Abstract.
Inspired by the work of Soma and Watari, we introduce a tree structure on the subsemimodules of the semigroup $\Gamma$ associated with an irreducible plane curve singularity $(C,O)$. Building on results of Oblomkov, Rasmussen, and Shende, we show that for certain classes of singularities this tree captures essential geometric information about the punctual Hilbert schemes of $(C,O)$. As an application, we derive an explicit formula for the motivic Hilbert zeta function of simple curve singularities.
A point of the Hilbert scheme corresponds to an ideal of the local ring $\mathcal{O}_{C,O}$ of the singularity. We investigate the stratification of these Hilbert schemes arising from conditions on the minimal number of generators of the defining ideals, and we describe geometric features of the resulting strata, including their dimension and closure relations. We then study a bivariate motivic zeta function that naturally extends a zeta function with integer coefficients and that, in the case of a plane curve singularity, has recently been related to the topology of the link of the singularity.
@@ 1. Introduction
From an algebraic viewpoint, let $A=\mathcal{O}_{C,O}$ denote the local ring of $C$ at $O$. Then the $k$-points of $C^{[\ell]}$ are in bijection with ideals of colength $\ell$ in $A$:
@@
Pfister and Steenbrink [undefo] proved that the punctual Hilbert schemes $C^{[l]}$ can be embedded in a closed linear subvariety of the Grassmannian $Gr(\delta,\overline{A}/I(2\delta)),$ where $I(2\delta)$ is the ideal consisting of elements $h\in\overline{A}$ (or equivalently of $A$) whose valuation $v(h)$ is larger than or equal to $2\delta.$ Moreover, the authors ”provided” the defining equations of this embedding. In general, even when the defining equations of punctual Hilbert schemes are known, it remains difficult to understand their geometry or to compute their invariants.
Progress in understanding the geometry of punctual Hilbert schemes has been relatively slow, due in part to the complexity of their structure even in simple singular cases. Soma and Watari [undeft, undefu] studied the geometry of punctual Hilbert schemes for plane curve singularities of types $A_{2d}$, $E_{6}$, and $E_{8}$. In the case of unibranched curve singularities, the punctual Hilbert scheme $C^{[\ell]}$ coincides with the compactified Jacobian $\overline{JC}$ for sufficiently large $\ell$ [undefp].
For curves defined by the equation $y^{m}-x^{n}=0$ with $\gcd(m,n)=1$, Lusztig and Smelt [undefk] computed the homology of the associated Hilbert schemes. Piontkowski [undefp] later extended these results to curves with a single Puiseux pair. Subsequently, Gorsky, Mazin, and Oblomkov [undefd] derived explicit formulas for the Poincaré polynomials of Hilbert schemes associated with singularities having Puiseux exponents $(nd,md,md+1)$.
A key insight in this line of research is that the intersection of the compactified Jacobian $\overline{JC}$ with certain Schubert cells in the affine Grassmannian has the structure of an affine space. This geometric fact admits rich combinatorial interpretations, particularly in relation to $(q,t)$-Catalan theory [undefe, undeff, undefg, undefi].
For curve singularities with a single Puiseux pair, Oblomkov, Rasmussen, and Shende [undefm] proved that the punctual Hilbert scheme $C^{[\ell]}$ admits a stratification by affine spaces, a result that plays a foundational role in the motivic study of these moduli spaces.
@@
On one hand, these schemes are the fundamental building blocks for the motivic Hilbert zeta function, introduced by Bejleri, Ranganathan and Vakil  [undef]. This zeta function is defined as the generating series of the classes of punctual Hilbert schemes in the Grothendieck ring of varieties:
@@
Compared to the classical motivic zeta function of Kapranov [undefj], the motivic Hilbert zeta function provides a finer invariant, capturing detailed information about the singularities of a variety $X$ defined over an algebraically closed field $k$.
On the other hand, a series of conjectures by Oblomkov, Rasmussen, and Shende [undefn, undefm] posits a deep connection between the geometry of punctual Hilbert schemes of plane curve singularities and the topology of associated algebraic links, when working over $\mathbb{C}$. More precisely, they conjecture that the (reduced) Khovanov–Rozansky homology of the link of the singularity can be recovered from the homology of the Hilbert schemes. Maulik [undefl] proved the first of these conjectures using an inductive argument based on blow-ups, analyzing the variation of both Hilbert schemes and the HOMFLY polynomial under such transformations.
Very recently, Rossinelli [undefr] discovered a new connection between curvilinear punctual Hilbert schemes and the Igusa motivic zeta function, further highlighting the role of Hilbert schemes as a unifying object in singularity theory and motivic integration. Diaconescu, Porta, Sala and Vosoughinia [undefb] made a progress in the case of space curve singularities. Fasola, Graffeo, Lewanski and Ricolfi [undefc] computed the motivic classes of punctual nested Hilbert schemes on a smooth pointed surface. Graffeo, Monavari, Moschetti, and Ricolfi [undefh] proved a closed formula for the generating function of the motives $\mathrm{Hilb}^{d}(\mathbb{A}^{n})_{0}]\in K_{0}(\mathrm{Var}_{\mathbb{C}})$ of punctual Hilbert schemes, summing over $n$, for fixed $d>0$.
@@
The latter class consists of singularities for which, within their equisingularity class—i.e., the set of curve singularities sharing the same semigroup—there exists a unique representative up to analytic isomorphism (see Section 6 for futher details). This class was originally introduced by Pfister and Steenbrink [undefo], and plays a distinguished role in the classification of unibranch curve singularities due to its rigidity under deformation.
To achieve this, we make extensive use of $\Gamma$-subsemimodules, where $\Gamma$ is the semigroup of an irreducible curve singularity. A $\Gamma$-subsemimodule is a subset $\Delta\subset\Gamma$ such that $\Gamma+\Delta\subseteq\Delta$. It follows directly from the properties of ideals and the valuation $v$ that, for any ideal $I\subset A$, the value set
@@
Inspired by recent work of Soma and Watari [undefu] on the geometry of Hilbert schemes, as well as that of Oblomkov, Rasmussen, and Shende [undefm] on knot invariants and curve singularities, we associate to a curve singularity with semigroup $\Gamma$ a leveled graph $G_{\Gamma}$ defined as follows.
@@
We study the graph $G_{\Gamma}$ and prove that it has the structure of a tree (Theorem 3.1): the important thing about this theorem is that it explains how to obtain all $\Gamma$-submodules. Now, each subsemimodule $\Delta\in\mathscr{D}_{\ell}$ defines a constructible subset $C^{[\Delta]}\subseteq C^{[\ell]}$, consisting of all ideals $I\subset A$ in $C^{[\ell]}$ such that $v(I\setminus\{0\})=\Delta$. One of the main result of the paper, is that, for certain curve singularities an edge of the graph induces a peicewise trivial fibration (the proof uses the work in [undefm] regarding the defining equations of $C^{[\Delta]}$); More precisely, we have the following theorem:
@@
We explicitly determine the image of this morphism. In the case of plane curve singularities defined by $x^{p}-y^{q}=0$ with $\gcd(p,q)=1$, or more generally for singularities with monomial semigroups, this allows us to recover a key result from [undefm]: $C^{[\Delta]}$ is an affine space, and its dimension is completely determined by the combinatorial invariants of the $\Gamma$-subsemimodule $\Delta$.
We also analyze an example of a plane curve singularity with two Puiseux pairs (see Example 4.3). In this case, the morphism
@@
still induces a piecewise trivial fibration, but the dimension of the fiber is not solely governed by the number of minimal generators below $\gamma_{\Delta}$. New phenomenas arise and must be further studied to fully understand the geometry of the strata and the structure of this morphism; we think that it is a fibration.
One important application of Theorem 1.1 is the development of an algorithm to compute the motivic Hilbert zeta function for irreducible curve singularities, as introduced in [undef]:
@@
For the singularities of types $A_{2d}$, $E_{6}$, and $E_{8}$, these formulas were independently obtained by Watari [undefv], using a different approach based on explicit geometric analysis and results from [undefu, undeft]. Our method, based on the tree structure $G_{\Gamma}$ and piecewise trivial fibrations, provides a unified and algorithmic derivation that extends naturally to more general singularities, including $W_{8}$ and $Z_{10}$.
In Section 5, in relation to the conjectures of Oblomkov, Rasmussen, Shende, and Maulik [undefn, undefm, undefl], we study the defining equations of certain stratifications of the punctual Hilbert schemes. For a given $\Gamma$-subsemimodule $\Delta$ and an integer $m\in\mathbb{N}$, we consider the
@@
where $n$ is the number of minimal generators of $\Delta$ as a $\Gamma$-subsemimodule, and $N(\Delta)$ is the dimension of $C^{[\Delta]}$ which is an affine space proven in [undefm, Theorem 13].
@@
This series $Z^{\mathrm{Hilb}}_{(C,O)}(a^{2},q^{2})$, introduced in this paper, naturally generalizes the zeta function studied in [undefn]. There, the Euler characteristic specialization (i.e., the point-counting or topological Euler characteristic) of this series was conjecturally related to the HOMFLY polynomial of the associated algebraic knot. Our construction lifts this invariant to the Grothendieck ring of varieties $K_{0}({\mathrm{Var}}_{{\mathbb{C}}})$, thereby encoding richer geometric information about the stratification of the Hilbert schemes by the number of generators.
@@ 2. Punctual Hilbert schemes
In this section, we recall the definition of the punctual Hilbert schemes of a curve singularity and review some of its fundamental properties. Let $(C,O)$ be the germ of a unibranch curve singularity defined over an algebraically closed field $k$ of characteristic zero, and let $A:=\mathcal{O}_{C,O}$ denote its local ring at $O$. Then $A$ is a one-dimensional local domain, and its integral closure $\overline{A}$ in its field of fractions is a discrete valuation ring, isomorphic to $k[[t]]$.
@@
and $c=2\delta$ if and only if $A$ is Gorenstein [undefs].
@@
Pfister and Steenbrink [undefo] showed that the punctual Hilbert schemes of a unibranch curve singularity can be embedded as closed subvarieties of a Grassmannian. More precisely, let $\mathscr{M}$ denote the reduced subscheme of the Grassmannian $\operatorname{Gr}(\delta,\overline{A}/I(2\delta))$ defined by
@@
Since $c\leq 2\delta$, this implies that for $\ell\geq 2\delta$, the structure of $C^{[\ell]}$ stabilizes in the sense that its points are completely determined by the fixed ambient variety $\mathscr{M}$. Consequently, to understand the classes $C^{[\ell]}]$ in the Grothendieck ring of varieties $K_{0}(\mathrm{Var}_{k})$, it suffices to study the punctual Hilbert schemes for $\ell\leq 2\delta$.
@@
[undefq, Lemme 5.1.24] For positive integer $\ell$, an ideal $I$ of $A$ belongs to $C^{[\ell]}$ if and only if $v(I\setminus\{0\})\in\mathscr{D}_{\ell}$.
@@
It is shown in [undefo, Lemma 5] that each stratum $C^{[\Delta]}$ arises as the intersection of $C^{[\ell]}$ with a Schubert cell in the Grassmannian $\operatorname{Gr}(\delta,\overline{A}/I(2\delta))$. In particular, this implies that $C^{[\Delta]}$ is a open subset of $C^{[\ell]}$, and its geometry is compatible with the cellular decomposition of the ambient Grassmannian.
@@ 3. The tree structure of subsemimodules
Let $\Gamma$ be the value semigroup associated with the germ of an irreducible curve singularity $(C,O)$. The goal of this section is to endow the set of $\Gamma$-subsemimodules with a tree structure. We will show in the next section that this combinatorial object plays a crucial role in understanding the geometry of the punctual Hilbert schemes of $C$. This construction is inspired by the work of Soma and Watari [undefu, undeft].
@@
The following result appears in [undefu]:
@@
The answer is affirmative, as will follow from Proposition 3.1, which shows that the deletion maps are jointly surjective. To prepare for this result, we first establish two lemmas.
@@
    Acyclicity: it follows from the uniqueness of Frobenius element of a $\Gamma$-subsemimodule.
@@ 4. Piecewise fibrations induced by the edges of the $\Gamma$-subsemimodules tree
Corollary 4.2 provides a new perspective on the geometry of $C^{[\Delta]}$ in the case $C=\{y^{p}=x^{q}\}$, recovering a result from [undefm, Theorem 13] via our combinatorial and valuative approach. Our method relies on the explicit defining equations of the strata $C^{[\Delta]}$, introduced in [undefm, Proposition 12].
We begin by recalling some key constructions from [undefm]. Let $(C,O)$ be the germ of a unibranch curve singularity—either of type $(p,q)$ or with a monomial semigroup—and let $A:=\mathcal{O}_{C,O}\subset{\mathbb{C}}[[t]]$ denote its complete local ring. We fix a basis of $A$ that is compatible with the monomial basis of the semigroup algebra ${\mathbb{C}}[\Gamma]$. Specifically, for each $i\in\Gamma$, we consider elements of the form
@@
$f_{\gamma_{j}}=\phi_{\gamma_{j}}+\sum_{k\in\Gamma_{>\gamma_{j}}\setminus\Delta}\lambda_{j}^{k-\gamma_{j}}\phi_{k}\in\mathcal{O}_{C}$
@@
$\tau_{\gamma_{j}}(\lambda^{\bullet}_{\bullet})=\phi_{\gamma_{j}}+\sum_{k\in\Gamma_{>\gamma_{j}}\setminus\Delta}\lambda_{j}^{k-\gamma_{j}}\phi_{k},\quad\lambda_{j}^{k-\gamma_{j}}\in{\mathbb{C}}[[\lambda^{\bullet}_{\bullet},t]].$
@@
This bijectivity follows from [undefn, Theorem 27], which ensures that every ideal in $C^{[\Delta]}$ admits a unique normalized generating set of the prescribed form. Consequently, the stratum $C^{[\Delta]}$ embeds into the parameter space $\operatorname{Gen}_{\Delta}$, yielding a locally closed embedding
@@
It is important to note that the inclusion $C^{[\Delta]}\hookrightarrow\operatorname{Gen}_{\Delta}$ does not imply that every point in $\operatorname{Gen}_{\Delta}$ corresponds to an ideal $J$ with $v(J\setminus\{0\})=\Delta$. This phenomenon is illustrated by the following example, adapted from [undefn]. Let $\mathcal{O}_{C}={\mathbb{C}}[[t^{3},t^{7}]]$, whose value semigroup is $\Gamma=\langle 3,7\rangle$. Consider the $\Gamma$-subsemimodule $\Delta=(6,10)_{\Gamma}$, minimally generated by $6$ and $10$. The element
@@
where $S$ denotes the composition ${\mathbb{C}}[\Gamma]^{\oplus m}\to\ker G\hookrightarrow{\mathbb{C}}[\Gamma]^{\oplus n}$. According to [undefp, Lemma 4], the kernel $\ker G$ is generated by homogeneous elements of the form
@@
$(s_{i})_{j}=u_{i}^{j}t^{\sigma_{i}-\gamma_{j}},\quad u_{i}^{j}\in{\mathbb{C}}.$
@@
1. We define the Augmented Syzygy of $\Delta$ as